# High-dimensional analysis of ridge regression for non-identically distributed data with a variance profile

### v2

@@ 2.1 High-dimensional linear regression from the random matrix perspective
When the sample size is comparable to the dimensionality of the observations, recent advances in RMT have been successfully applied to various inference problems in high-dimensional multivariate statistics, see e.g. [NPW21] for a recent overview. Many works have considered the high-dimensional analysis of the linear model using tools from RMT for iid data with a general covariance structure $\Sigma\in{\mathbb{R}}^{p\times p}$ (assumed to be a positive semi-definite matrix) that is for
[…]
In particular, for such data, the study of the minimum norm least-square estimator and the double descent behavior of the predictive risk has been considered in [HMRT22, Bac24, BHX20, RMR21]. The analysis of the predictive risk of ridge regression using iid data with a general covariance structure has been studied in [DW18, Bac24], while previous works on the statistical analysis of ridge regression from the RMT perspective include [EK18, Dic16] and [CD11, TV04] for applications in wireless communication.
@@ 2.2 Linear regression for independent but non-identically distributed data
While the statistical analysis of linear regression for iid data with a general covariance structure is very well understood, the literature on the study of the linear model for non-identically distributed predictors appears to be scarcer. A first analysis of maximum likelihood estimation in standard models (including linear regression) for independent but non-identically distributed data dates back to [Ber82]. More recent works [BBK+19, KBB+20], on statistical inference in linear regression in the so-called model-free framework, allow to consider the setting on non-identically distributed predictors.
@@ 2.3 The use of variance profile in RMT
RMT allows to describe the asymptotic distribution of the eigenvalues of large matrices with random entries, see e.g. [BS10]. In particular, the well-known Marchenko-Pastur theorem characterizes the limiting spectral distribution of the covariance matrix $\widehat{\Sigma}_{n}=\frac{1}{n}X_{n}^{t}X_{n}$ for a data matrix $X_{n}=Z_{n}\Sigma^{1/2}$ with iid rows in the asymptotic setting $\lim\limits_{n\to+\infty}\frac{p}{n}=c>0$.
However, in many applications, such as photon imaging [SHDW14], network traffic analysis [BMG13], ecology [AGB+15, AT15], neurosciences [ASS15, ARS15] or genomics for microbiome studies [CZL19], the amount of variance in the observed data matrix may significantly change from one sample to the other, that is between the rows of $X_{n}$. The literature on statistical inference from high-dimension matrices with heteroscedasticity has thus recently been growing [BM20, BDF17, LDS18, UHZB16, ZCW22, JFL22]. Modeling data as a random matrix $X_{n}=\Upsilon_{n}\circ Z_{n}$ with a variance profile to handle the setting of non-iid data has also found applications in the analysis of the performances of wireless digital communication channels [HLN07]. In the RMT literature, Hermitian random matrices with centered entries but non-equal distribution are referred to as generalized Wigner matrices for which many asymptotic properties are now well understood. For example, for Hermitian random matrices with a variance profile that is doubly stochastic (namely its rows and columns elements sump up to one), bulk universality at optimal spectral resolution for local spectral statistics have been established in [EYY12] and they are shown to converge to those of a standard Wigner matrix (that is with iid sub-diagonal entries). The case of a generalized Wigner matrix with a variance profile that is not necessarily doubly stochastic has been studied in [AEK17a], and non-herminitian random matrices with a variance profile have been considered in [CHNR18, HLN07, HLN06] using the notion of deterministic equivalent that consists in approximating the spectral distribution of a random matrix by a deterministic function. A recent work [BHX23] considers the analysis of approximate message passing for statistical estimation problems involving a data matrix with a variance profile, and an application to ridge regression to characterize the non-asymptotic distribution of the ridge estimator is proposed.
@@
Let $\mu$ be a probability measure supported on ${\mathbb{R}}$. Then, its Stieltjes transform is defined as
@@
where +=$\mathbb{C}^{+}=\{z\in\mathbb{C},\Im(z)>0\}$ and $\Im(\cdot)$ denotes the imaginary part of a complex number.
@@
of the empirical eigenvalue distribution $\hat{\mu}_{n}$ of $\widehat{\Sigma}_{n}$, and the empirical eigenvalue distribution $\tilde{\mu}_{n}$ of $\widetilde{\Sigma}_{n}=\frac{1}{n}X_{n}X_{n}^{t}$ respectively.
To obtain these deterministic equivalents, the following assumptions will be needed, on mild moment conditions on the $Z_{ij}$’s and boundedness hypotheses on the entries of the variance profile matrix.
Assumption 2.1.
There exists $\delta>0$ such that, $\mathbb{E}[|Z_{ij}|^{4+\delta}]<+\infty,$ for all $1\leq i\leq n$ and $1\leq j\leq p$.
Assumption 2.2.
There exists $\gamma_{\mathrm{max}}>0$ such that, $\underset{n\geq 1}{\mathrm{sup}}\underset{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq p\end{subarray}}{\mathrm{max}}|\gamma_{ij}^{(n)}|<\gamma_{\mathrm{max}}.$
Assumption 2.3.
There exists $\gamma_{\mathrm{min}}>0$ such that, $\forall n\geq 1,\underset{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq n\end{subarray}}{\mathrm{min}}|\gamma_{ij}^{(n)}|\geq\gamma_{\mathrm{min}}.$
Assumption 2.4.
There exists $d_{\ast}>0$ such that, $|\frac{p}{n}-1|\geq d_{\ast}$ for all values of $n$ and $p$.
Throughout the paper, we also assume that $\widetilde{X}=\widetilde{S}_{p}^{1/2}\widetilde{Z}$ with
$\mathbb{E}[|\widetilde{Z}_{j}|^{4+\delta}]<+\infty,\quad\underset{\begin{subarray}{c}1\leq j\leq p\end{subarray}}{\mathrm{max}}|\tilde{\gamma}_{j}^{(n)}|\leq\gamma_{\mathrm{max}}\;\mbox{and}\;\underset{\begin{subarray}{c}1\leq j\leq n\end{subarray}}{\mathrm{min}}|\tilde{\gamma}_{j}^{(n)}|\geq\gamma_{\mathrm{min}}$
and for all $n\geq 1$. Then, using [HLN07][Theorem 2.5], the following result holds.
Proposition 2.1.
Under Assumptions 2.1-2.4, the following limit holds true almost surely
@@
$\frac{1}{T_{p}(z)}=-zI_{p}+\frac{1}{n}\Gamma_{n}^{t}\frac{1}{1+\frac{1}{n}\Gamma_{n}T_{p}(z)},$ (2.3)
@@
$\frac{1}{\widetilde{T}_{n}(z)}=-zI_{n}+\frac{1}{n}\Gamma_{n}\frac{1}{1+\frac{1}{n}\Gamma_{n}^{t}\widetilde{T}_{n}(z)}.$
In this paper, we sometimes consider, as an illustrative example, a specific class of variance profiles said to be quasi doubly stochastic in the following sense.
Definition 2.2.
The variance profile $\Gamma_{n}$ is called quasi doubly stochastic when its rows and columns sum up to constant values (that are necessarily different when $n\neq p$), namely
$\frac{1}{n}\sum_{j=1}^{p}\gamma_{ij}^{2}=\frac{p}{n},\mbox{ for all }1\leq i\leq n,\quad\mbox{and}\quad\frac{1}{n}\sum_{i=1}^{n}\gamma_{ij}^{2}=1,\mbox{ for all }1\leq j\leq p.$
The fixed point equation (2.3) typically does not have an explicit expression, and one has to rely on numerical methods to solve it as done in Section 4. Nevertheless, when the variance profile is quasi doubly stochastic, the solution of the Dyson equation is a vector $T_{p}(z)$ having constant entries equal to $m_{p}(z)\in\mathbb{C}$ (or equivalently $T_{p}(z)=m_{p}(z)I_{p}$ is a scalar matrix) that satisfies
@@ 3 Main results
In this section, we derive deterministic equivalents for the DOF and the predictive risk of ridge regression. We also obtain a deterministic equivalent of the predictive risk of minimum norm least square estimation when the ridge regularization parameter tends to zero. We compare these results to those that are already known in the standard setting of iid data, and we highlight the emergence of the double descent phenomenon for non-iid data. Throughout this section, it is supposed that Assumptions 2.1-2.4 hold true.
@@ 3.1 Degrees of freedom
For any $\lambda>0$, the following holds
@@
$\kappa(\lambda)$ $=$ $\mathop{\rm diag}\nolimits(\kappa_{1}(\lambda),\ldots,\kappa_{p}(\lambda)),\quad\mbox{where}\quad\kappa_{j}(\lambda)=\frac{\operatorname{Tr}[\widetilde{D}_{j}]}{\operatorname{Tr}[\widetilde{D}_{j}\widetilde{T}_{n}(-\lambda)]},\quad\mbox{for}\quad 1\leq j\leq p.$ (3.2)
When the variance profile is quasi doubly stochastic, one has that $\operatorname{Tr}[\widetilde{D}_{j}]=n$. Moreover, as remarked in Section 2.3, the matrix $\widetilde{T}_{n}(-\lambda)=\tilde{m}_{n}(-\lambda)I_{n}$ is scalar, implying that
$\operatorname{Tr}[\widetilde{D}_{j}\widetilde{T}_{n}(-\lambda)]=n\tilde{m}_{n}(-\lambda),\quad\mbox{for all}\quad 1\leq j\leq p.$
Moreover, in this setting, $\Sigma_{n}=I_{p}$. Consequently, if $\Gamma_{n}$ is quasi doubly stochastic, one has that
$\kappa(\lambda)=\frac{1}{\tilde{m}_{n}(-\lambda)}I_{p},$
and the deterministic equivalent of the DOF is
@@
3.2 Predictive risk
We first express the predictive risk in a more convenient way. To this end, for any square matrix $A$, we denote by $\Delta[A]$ the diagonal matrix whose diagonal entries are those of $A$.
@@ 3.2 Predictive risk
For any $\lambda>0$, the predictive risk $\hat{r}_{\lambda}(X_{n})=\mathbb{E}[(\widetilde{Y}-\widetilde{X}^{t}\hat{\theta}_{\lambda})^{2}|X_{n}]$ has the following expression
$\hat{r}_{\lambda}(X_{n})$ $=$ $\sigma^{2}+\frac{\sigma^{2}}{n}\mathrm{Tr}\big{[}\widetilde{S}_{p}\Delta[Q_{p}(-\lambda)]\big{]}+\lambda\left(\frac{\lambda\alpha^{2}}{p}-\frac{\sigma^{2}}{n}\right)\mathrm{Tr}\big{[}\widetilde{S}_{p}\Delta[Q_{p}^{\prime}(-\lambda)]\big{]}.$ (3.3)
Moreover, we can exhibit the following Bias-Variance decomposition for $\hat{r}_{\lambda}(X_{n})$ :
$\hat{r}_{\lambda}(X_{n})$ $=$ $\sigma^{2}+\mathrm{Bias}(\hat{\theta}_{\lambda})+\mathrm{Var}(\hat{\theta}_{\lambda}),$
@@
$\mathop{\rm Bias}\nolimits(\hat{\theta}_{\lambda})$ $=$ $\mathbb{E}\big{[}(\mathbb{E}[\hat{\theta}_{\lambda}|X_{n},\theta_{\ast}]-\theta_{\ast})^{t}\widetilde{S}_{p}(\mathbb{E}[\hat{\theta}_{\lambda}|X_{n},\theta_{\ast}]-\theta_{\ast})\big{|}X_{n}\big{]}$
@@
$\mathop{\rm Var}(\hat{\theta}_{\lambda})$ $=$ $\mathbb{E}\big{[}(\hat{\theta}_{\lambda}-\mathbb{E}[\hat{\theta}_{\lambda}|X_{n},\theta_{\ast}])^{t}\widetilde{S}_{p}(\hat{\theta}_{\lambda}-\mathbb{E}[\hat{\theta}_{\lambda}|X_{n},\theta_{\ast}])\big{|}X_{n}\big{]}$
$=$ $\frac{\sigma^{2}}{n}\mathrm{Tr}\big{[}\widetilde{S}_{p}\Delta[Q_{p}(-\lambda)]-\lambda\widetilde{S}_{p}\Delta[Q_{p}^{\prime}(-\lambda)]\big{]},$
where
$Q_{p}(z)=(\widehat{\Sigma}_{n}-zI_{p})^{-1}\quad\mbox{for}\quad z\in\mathbb{C}\setminus{\mathbb{R}}^{+},$
is the resolvent of the matrix $\widehat{\Sigma}_{n}$, and $Q_{p}^{\prime}$ denotes the derivative of $Q_{p}$ with respect to $z$.
We can now give a deterministic equivalent of the predictive risk $\hat{r}_{\lambda}(X_{n})$ that is obtained in a simple way by replacing the diagonal matrix $\Delta[Q_{p}(z)]$ with the deterministic matrix $T_{p}(z)$ in the expression (3.3) of the predictive risk. All proofs of the following results are given in Appendix A.3.
@@
A deterministic equivalent of the predictive risk is
$r_{\lambda}(\Upsilon_{n})=\sigma^{2}+\frac{\sigma^{2}}{n}\mathrm{Tr}[\widetilde{S}_{p}T_{p}(-\lambda)]+\lambda\left(\frac{\lambda\alpha^{2}}{p}-\frac{\sigma^{2}}{n}\right)\mathrm{Tr}[\widetilde{S}_{p}T_{p}^{\prime}(-\lambda)],$ (3.4)
@@
$\lim_{n\to\infty,\;p/n\to c}|\hat{r}_{\lambda}(X_{n})-r_{\lambda}(\Upsilon_{n})|=0,\;\mbox{almost surely},$
where $T_{p}^{\prime}(z)$ denotes the derivative of $T_{p}(z)$ with respect to $z$.
A key argument to derive the computation of $r_{\lambda}(\Upsilon_{n})$ is to prove that the matrix $T_{p}(z)$ is a relevant deterministic equivalent of the diagonal matrix $\Delta[Q_{p}(z)]$ for an appropriate notion of asymptotic equivalence between matrices of growing size. This is the purpose of Apppendix A.2 where we derive a stronger convergence result than the one stated in Proposition 2.1 which only shows that $\frac{1}{p}\operatorname{Tr}[T_{p}(z)]$ is a deterministic equivalent of $g_{\hat{\mu}_{n}}(z)=\frac{1}{p}\operatorname{Tr}[Q_{p}(z)]$.
@@
If $\frac{p}{n}<1$ then $T_{p}(-\lambda)$ and its derivative, $T_{p}^{\prime}(-\lambda)$ admit a limit when $\lambda$ tends to 00. Indeed, there exists a constant $\tau>0$ such that
@@
Suppose that the variance profile $\Gamma_{n}$ is quasi doubly stochastic. Then, if $\frac{p}{n}<1$, one has that
$T_{p}(0^{-})=m_{p}(0)I_{p},$
and, if $\frac{p}{n}>1$,
$\kappa(0^{+})=\frac{1}{\tilde{m}_{n}(0)}I_{p}\quad\mathrm{and}\quad\kappa^{\prime}(0^{+})=\frac{\tilde{m}_{n}^{\prime}(0)}{\tilde{m}_{n}^{2}(0)}I_{p},$
where $m_{p}(.)$, resp. $\tilde{m}_{n}(.)$, is the Stieltjes transform of the Marchenko-Pastur distribution with parameter $p/n$, resp. $n/p$.
Then, Theorem 3.1 allows us to understand the behavior of $r_{\lambda}(\Upsilon_{n})$ when $\lambda\rightarrow 0$ and $\lambda\rightarrow+\infty$ through the following corollary.

### v3

@@ 2.1 High-dimensional linear regression from the random matrix perspective
When the sample size is comparable to the dimensionality of the observations, recent advances in RMT have been successfully applied to various inference problems in high-dimensional multivariate statistics, see e.g. [NPW21] for a recent overview. Many works have considered the high-dimensional analysis of the linear model using tools from RMT for iid data with a general covariance structure $\Sigma\in{\mathbb{R}}^{p\times p}$ (assumed to be a positive semi-definite matrix) that is for $X_{n}=Z_{n}\Sigma^{1/2}$, for an $n\times p$ matrix $Z_{n}$ with iid centered entries having variance one.
In particular, for such data, the study of the minimum norm least-square estimator and the double descent behavior of the predictive risk has been considered in [HMRT22, Bac24, BHX20, RMR21]. The analysis of the predictive risk of ridge regression using iid data with a general covariance structure has been studied in [DW18, Bac24], while previous works on the statistical analysis of ridge regression from the RMT perspective include [EK18, hD16] and [CD11, TV04] for applications in wireless communication.
@@ 2.2 Linear regression for independent but non-identically distributed data
While the statistical analysis of linear regression for iid data with a general covariance structure is very well understood, the literature on the study of the linear model for non-identically distributed predictors appears to be scarcer. A first analysis of maximum likelihood estimation in standard models (including linear regression) for independent but non-identically distributed data dates back to [Ber82]. More recent works [BBK+19, KBB+20], on statistical inference in linear regression in the so-called model-free framework, allow to consider the setting on non-identically distributed predictors. The assumption of independent and identically distributed (iid) data often fails in real-world scenarios. This limitation notably appears in fields such as epidemiology [ETP+22], finance [Gro21], neuroscience [TvWM+13], and climatology [ZNM+21]. The recognition of this issue has motivated significant research into the analysis of non-iid data. For example, [ZMP22] highlights the non-iid nature of data in studies regarding the effectiveness of COVID-19 vaccines. They demonstrate that interactions between data points can introduce bias into estimation results. Beyond the risk of bias, the non-iid structure of data can also invalidate certain statistical methods. Generalized cross-validation, for instance, yields poor predictions in such contexts. In response to this challenge, [LLS24, AZVP24] have proposed robust alternatives specifically designed to handle dependencies between data points. The study of dependent data has also garnered growing attention in other domains. In control theory, researchers have made strides in addressing dependencies [NWB+20, TZMP23, ZTPM24]. Similarly, the field of signal denoising has seen advancements with methods that explicitly account for data correlations [SN23, KSS23]. These developments underscore the importance of adapting traditional methodologies to the challenges posed by non-iid data across a broad spectrum of disciplines. The study of correlated data is a growing topic in high-dimensional statistics [ZJVM24, ZM24]. However, to the best of our knowledge, the high-dimensional analysis of the linear model using non-identically distributed data has not considered so far. Hence we propose to tackle this framework by endowing variance profile to the data.
@@ 2.3 The use of variance profile in RMT
RMT allows to describe the asymptotic distribution of the eigenvalues of large matrices with random entries, see e.g. [BS10]. In particular, the well-known Marchenko-Pastur theorem characterizes the limiting spectral distribution of the covariance matrix $\widehat{\Sigma}_{n}=\frac{1}{n}X_{n}^{\top}X_{n}$ for a data matrix $X_{n}=Z_{n}\Sigma^{1/2}$ with iid rows in the asymptotic setting $\lim\limits_{n\to+\infty}\frac{p}{n}=c>0$.
However, in many applications, such as photon imaging [SHDW14], network traffic analysis [BMG13], ecology [AGB+15, AT15], neurosciences [ASS15, ARS15] or genomics for microbiome studies [CZL19], the amount of variance in the observed data matrix may significantly change from one sample to the other, that is between the rows of $X_{n}$. The literature on statistical inference from high-dimension matrices with heteroscedasticity has thus recently been growing [BM20, BDF17, LDS18, UHZB16, ZCW22, JFL22]. Modeling data as a random matrix $X_{n}=\Upsilon_{n}\circ Z_{n}$ with a variance profile to handle the setting of non-iid data has also found applications in the analysis of the performances of wireless digital communication channels [HLN07]. In the RMT literature, Hermitian random matrices with centered entries but non-equal distribution are referred to as generalized Wigner matrices for which many asymptotic properties are now well understood. For example, for Hermitian random matrices with a variance profile that is doubly stochastic (namely its rows and columns elements sump up to one), bulk universality at optimal spectral resolution for local spectral statistics have been established in [EYY12] and they are shown to converge to those of a standard Wigner matrix (that is with iid sub-diagonal entries). The case of a generalized Wigner matrix with a variance profile that is not necessarily doubly stochastic has been studied in [AEK17a], and non-hermitian random matrices with a variance profile have been considered in [CHNR18, HLN07, HLN06] using the notion of deterministic equivalent that consists in approximating the spectral distribution of a random matrix by a deterministic function. A recent work [BHX23] considers the analysis of approximate message passing for statistical estimation problems involving a data matrix with a variance profile, and an application to ridge regression to characterize the non-asymptotic distribution of the ridge estimator is proposed.
@@
of the empirical eigenvalue distribution $\hat{\mu}_{n}$ of $\widehat{\Sigma}_{n}$, and the empirical eigenvalue distribution $\tilde{\mu}_{n}$ of $\widetilde{\Sigma}_{n}=\frac{1}{n}X_{n}X_{n}^{\top}$ respectively. Then, using [HLN07][Theorem 2.5], the following result holds.
Theorem 2.1.
Under Assumptions 1.2-1.5, the following limit holds true almost surely
@@
$\frac{1}{T_{p}(z)}=-zI_{p}+\frac{1}{n}\Gamma_{n}^{\top}\frac{1}{1+\frac{1}{n}\Gamma_{n}T_{p}(z)},$ (2.3)
@@
$\frac{1}{\widetilde{T}_{n}(z)}=-zI_{n}+\frac{1}{n}\Gamma_{n}\frac{1}{1+\frac{1}{n}\Gamma_{n}^{\top}\widetilde{T}_{n}(z)}.$
In this paper, we sometimes consider, as an illustrative example, the specific class of quasi doubly stochastic variance profiles defined in Equation 1.6. The fixed point equation (2.3) typically does not have an explicit expression, and one has to rely on numerical methods to solve it as done in Section 4. Nevertheless, when the variance profile is quasi doubly stochastic, the solution of the Dyson equation is a vector $T_{p}(z)$ having constant entries equal to $m_{p}(z)\in\mathbb{C}$ (or equivalently $T_{p}(z)=m_{p}(z)I_{p}$ is a scalar matrix) that satisfies
@@ 3 Main results
In this section, we derive deterministic equivalents for the DOF and the predictive risk of ridge regression. We also obtain a deterministic equivalent of the predictive risk of minimum norm least square estimation when the ridge regularization parameter tends to zero. We compare these results to those that are already known in the standard setting of iid data, and we highlight the emergence of the double descent phenomenon for non-iid data. Throughout this section, it is supposed that Assumptions 1.2-1.5 hold true.
@@ 3.1 Degrees of freedom
Under Assumptions 1.2-1.3, for any $\lambda>0$, the following holds
@@
$\kappa(\lambda)$ $=$ $\mathop{\rm diag}\nolimits(\kappa_{1}(\lambda),\ldots,\kappa_{p}(\lambda)),\mbox{ where }\kappa_{j}(\lambda)=\frac{\operatorname{Tr}[\widetilde{D}_{j}]}{\operatorname{Tr}[\widetilde{D}_{j}\widetilde{T}_{n}(-\lambda)]}.$ (3.2)
When the variance profile is quasi doubly stochastic, one has that $\operatorname{Tr}[\widetilde{D}_{j}]=n$. Moreover, as remarked in Section 2.3, the matrix $\widetilde{T}_{n}(-\lambda)=\tilde{m}_{n}(-\lambda)I_{n}$ is scalar, implying that for $1\leq j\leq p$, $\operatorname{Tr}[\widetilde{D}_{j}\widetilde{T}_{n}(-\lambda)]=n\tilde{m}_{n}(-\lambda)$. Moreover, in this setting, $\Sigma_{n}=I_{p}$. Consequently, if $\Gamma_{n}$ is quasi doubly stochastic, one has that $\kappa(\lambda)=\frac{1}{\tilde{m}_{n}(-\lambda)}I_{p},$ and the deterministic equivalent of the DOF is
@@
3.2 Deterministic equivalents of the diagonal of the resolvent
Let $Q_{p}(z)=(\widehat{\Sigma}_{n}-zI_{p})^{-1}$ for $z\in\mathbb{C}\setminus{\mathbb{R}}^{+},$ be the resolvent of the matrix $\widehat{\Sigma}_{n}=\frac{1}{n}X_{n}^{\top}X_{n}$, and for any square matrix $A$, we denote by $\Delta[A]$ the diagonal matrix whose diagonal entries are those of $A$.
Then, as shown in the next subsection, a key argument to derive deterministic equivalents of the predictive and training risks is to prove that the matrix $T_{p}(z)$, defined by (2.3), is a relevant deterministic equivalent of the diagonal matrix $\Delta[Q_{p}(z)]$ for an appropriate notion of asymptotic equivalence between matrices of growing size. This is the purpose of Theorem 3.1 below that derives a stronger convergence result than the one stated in Proposition 2.1 which only shows that $\frac{1}{p}\operatorname{Tr}[T_{p}(z)]$ is a deterministic equivalent of $g_{\hat{\mu}_{n}}(z)=\frac{1}{p}\operatorname{Tr}[Q_{p}(z)]$. To this end, we define the following equivalence relation in order to specify an appropriate notion of deterministic equivalent for the matrix $\Delta[Q_{p}(-\lambda)]$.
Definition 3.1.
for all family of deterministic matrices $(U_{p})_{p\geq 1}$ satisfying:
$\forall p\geq 1,\ \ \ U_{p}\in\mathbb{R}^{p\times p}$ and $\exists K>0,\ \sup_{p\geq 1}\max_{i\in[p]}|U_{p}^{(i)}|\leq K,$ (3.3)
where $U_{p}^{(i)}$ denotes the i-th diagonal entry of $U_{p}$.
Then using this definition, we extend Theorem 2.1 from [HLN07] by proving that $T_{p}(z)$ is a relevant deterministic equivalent of $\Delta[Q_{p}(z)]$ in the following sense.
Theorem 3.1.
Define the following matrix families $\Delta[\mathbf{Q}(z)]=(\Delta[Q_{p}(z)])_{p\geq 1}$, $\Delta[\mathbf{Q^{\prime}}(z)]=(\Delta[Q^{\prime}_{p}(z)])_{p\geq 1}$, $\mathbf{T}(z)=(T_{p}(z))_{p\geq 1}$ and $\mathbf{T^{\prime}}(z)=(T^{\prime}_{p}(z))_{p\geq 1}$. Then, under Assumptions 1.2-1.3 the following equivalences hold:
We postpone the proof of this theorem to Appendix A.2 that is inspired from the proof of Theorem 2.1 in [HLN07]. This theorem is a cornerstone in the derivation of deterministic equivalents for $\hat{r}_{\lambda}^{test}(X_{n})$ and $\hat{r}_{\lambda}^{train}(X_{n})$ that leads to the main results of this paper provided in the next subsection.
3.3 Deterministic equivalents of the training and predictive risks
We first express the training and predictive risks in a more convenient way.
@@ 3.3 Deterministic equivalents of the training and predictive risks
For any $\lambda>0$, the training risk $\hat{r}_{\lambda}^{train}(X_{n})=\frac{1}{n}\mathbb{E}[\parallel Y_{n}-X_{n}\hat{\theta}_{\lambda}\parallel^{2}|X_{n}]$ and the predictive risk $\hat{r}_{\lambda}^{test}(X_{n})=\mathbb{E}[(\tilde{y}-\tilde{x}^{\top}\hat{\theta}_{\lambda})^{2}|X_{n}]$ have the following expressions
$\hat{r}_{\lambda}^{train}(X_{n})$ $=$ $\frac{\lambda^{2}\alpha^{2}}{n}\mathrm{Tr}\Bigg{[}Q_{p}(-\lambda)-\lambda Q_{p}^{\prime}(-\lambda)\Bigg{]}+\frac{\lambda^{2}\sigma^{2}}{n}\mathrm{Tr}[Q_{p}^{\prime}(-\lambda)],$ (3.4)
$\hat{r}_{\lambda}^{test}(X_{n})=\sigma^{2}+\frac{\sigma^{2}}{n}\mathrm{Tr}\big{[}\widetilde{S}_{p}\Delta[Q_{p}(-\lambda)]\big{]}+\lambda\left(\frac{\lambda\alpha^{2}}{p}-\frac{\sigma^{2}}{n}\right)\mathrm{Tr}\big{[}\widetilde{S}_{p}\Delta[Q_{p}^{\prime}(-\lambda)]\big{]},$ (3.5)
where $Q_{p}(z)=(\widehat{\Sigma}_{n}-zI_{p})^{-1}$ for $z\in\mathbb{C}\setminus{\mathbb{R}}^{+},$ is the resolvent of the matrix $\widehat{\Sigma}_{n}=\frac{1}{n}X_{n}^{\top}X_{n}$, and $Q_{p}^{\prime}(z)$ denotes the derivative of $Q_{p}(z)$ with respect to $z$. Moreover, we can exhibit the following Bias-Variance decomposition for $\hat{r}_{\lambda}^{test}(X_{n})$:
$\hat{r}_{\lambda}^{test}(X_{n})$ $=$ $\sigma^{2}+\mathrm{Bias}(\hat{\theta}_{\lambda})+\mathrm{Var}(\hat{\theta}_{\lambda}),$
@@
$\mathop{\rm Bias}\nolimits(\hat{\theta}_{\lambda})$ $=$ $\mathbb{E}\big{[}(\mathbb{E}[\hat{\theta}_{\lambda}|X_{n},\beta_{\ast}]-\beta_{\ast})^{\top}\widetilde{S}_{p}(\mathbb{E}[\hat{\theta}_{\lambda}|X_{n},\beta_{\ast}]-\beta_{\ast})\big{|}X_{n}\big{]}$
@@
$\mathop{\rm Var}(\hat{\theta}_{\lambda})$ $=$ $\mathbb{E}\big{[}(\hat{\theta}_{\lambda}-\mathbb{E}[\hat{\theta}_{\lambda}|X_{n},\beta_{\ast}])^{\top}\widetilde{S}_{p}(\hat{\theta}_{\lambda}-\mathbb{E}[\hat{\theta}_{\lambda}|X_{n},\beta_{\ast}])\big{|}X_{n}\big{]}$
$=$ $\frac{\sigma^{2}}{n}\mathrm{Tr}\big{[}\widetilde{S}_{p}\Delta[Q_{p}(-\lambda)]-\lambda\widetilde{S}_{p}\Delta[Q_{p}^{\prime}(-\lambda)]\big{]}.$
We can now give a deterministic equivalent of the predictive risk $\hat{r}_{\lambda}^{test}(X_{n})$ that is obtained in a simple way by replacing the diagonal matrix $\Delta[Q_{p}(z)]$ with the deterministic matrix $T_{p}(z)$ in the expression (3.5) of the predictive risk. All proofs of the following results are given in Appendix A.3.
@@
Under Assumptions 1.1 to 1.3, one can provide a deterministic equivalent for the predictive risk (respectively for the training risk), denoted by $r_{\lambda}^{test}$ (respectively $r_{\lambda}^{train}$) and defined as follows
$r_{\lambda}^{test}(X_{n})$ $=$ $\sigma^{2}+\frac{\sigma^{2}}{n}\mathrm{Tr}[\widetilde{S}_{p}T_{p}(-\lambda)]+\lambda\left(\frac{\lambda\alpha^{2}}{p}-\frac{\sigma^{2}}{n}\right)\mathrm{Tr}[\widetilde{S}_{p}T_{p}^{\prime}(-\lambda)],$ (3.6)
$r_{\lambda}^{train}(X_{n})$ $=$ $\frac{\lambda^{2}\alpha^{2}}{n}\mathrm{Tr}\Bigg{[}T_{p}(-\lambda)-\lambda T_{p}^{\prime}(-\lambda)\Bigg{]}+\frac{\lambda^{2}\sigma^{2}}{n}\mathrm{Tr}[T_{p}^{\prime}(-\lambda)],$ (3.7)
@@
$\lim_{n\to\infty,\;p/n\to c}|\hat{r}_{\lambda}^{\bullet}(X_{n})-r_{\lambda}^{\bullet}(X_{n})|=0,\;\mbox{almost surely},$
where $\bullet\in\{train,\ test\}$ and $T_{p}^{\prime}(z)$ denotes the derivative of $T_{p}(z)$ with respect to $z$.
From Theorem 3.2, one can deduce that, in the high dimensional regime, the training and predictive risk crystallize around deterministic values that depend on the data only through their variance profile. These deterministic equivalents can be expressed explicitly whenever the variance profile is quasi doubly-stochastic since $T_{p}(z)$ is equal to the Marchenko-Pastur Stieltjes transform in this case (see Section 2.3). However, the equivalents are not explicit in the general case since one does not have an explicit formula for $T_{p}(z)$. Nevertheless, $T_{p}(z)$ being the solution of a fixed-point equation, it can be approximated through a fixed-point algoritheorem. This alternative allowed us to perform numerical experiment that testify the accuracy of these deterministic equivalents (see Section 4).
@@
Under Assumptions 1.2 to 1.5, if $\frac{p}{n}<1$ then $T_{p}(-\lambda)$ and its derivative, $T_{p}^{\prime}(-\lambda)$ admit a limit when $\lambda$ tends to 00. Indeed, there exists a constant $\tau>0$ such that
@@
This lemma asserts that if $p<n$ (respectively $p>n$) then $T_{p}(-\lambda)$ and $T^{\prime}_{p}(-\lambda)$ (respectively $\kappa(\lambda)$) admit limits whenever $\lambda$ goes to 00. These limits ensures that the deterministic equivalents provided by Theorem 3.2 exist in the ridge (less) case, that is $\lambda=0$ (See Corollary 3.2).
@@
Suppose Assumptions 1.2-1.5 and that the variance profile $\Gamma_{n}$ is quasi doubly stochastic. Then, if $\frac{p}{n}<1$, one has that $T_{p}(0^{-})=m_{p}(0)I_{p},$ and, if $\frac{p}{n}>1$, $\kappa(0^{+})=\frac{1}{\tilde{m}_{n}(0)}I_{p}\quad\mathrm{and}\quad\kappa^{\prime}(0^{+})=\frac{\tilde{m}_{n}^{\prime}(0)}{\tilde{m}_{n}^{2}(0)}I_{p},$ where $m_{p}(.)$, resp. $\tilde{m}_{n}(.)$, is the Stieltjes transform of the Marchenko-Pastur distribution with parameter $p/n$, resp. $n/p$.
Then, Theorem 3.2 allows us to understand the behavior of $r_{\lambda}^{test}(X_{n})$ when $\lambda\rightarrow 0$ and $\lambda\rightarrow+\infty$ through the following corollary.